Found this tip:

To convert inches to millimeters. add 1/4 inch for every 15 inches, then multiply by 25. Accurate to within + - 0.09%
15 + 1/4 = 15.25 * 25 = 381.25
15.25 * 25.4 = 387.35
Where am I correct?

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Found this tip:

To convert inches to millimeters. add 1/4 inch for every 15 inches, then multiply by 25. Accurate to within + - 0.09%
15 + 1/4 = 15.25 * 25 = 381.25
15.25 * 25.4 = 387.35
Where am I correct?
==================
Where are you correct? You're correct if you get the right answer.
You're correct in Alaska, if you get the right answer.
You're correct in Alabama, if you get the right answer.
You're correct in Australia, if you get the right answer.
"Where am I correct?" makes no sense.
==============
1 inch = 25.4 mm
All you need to know.

Answer by ikleyn(52776)

.

The  "HINT"  you found is

        a)  incorrect;

        b)  is  NONSENSE.


1 inch = 25.4 millimeters.


All you need to know.


            P-E-R-I-O-D.